TPW, a calendar year taxpayer, sold land with a $543,000 tax basis for $790,000 in February. The purchaser paid $83,000 cash at closing and gave TPW an interest-bearing note for the $707,000 remaining price. In August, TPW received a $58,550 payment from the purchaser consisting of a $35,350 principal payment and a $23,200 interest payment. Assume that TPW uses the installment sale method of accounting.

  1. Compute the difference between TPW’s book and tax income resulting from the installment sale method.
  2. Is this difference favorable or unfavorable?
  3. Using a 21 percent tax rate, compute PTR’s deferred tax asset or liability (identify which) resulting from the book/tax difference.
  4. TPW, a calendar year taxpayer, sold land with a $543,000 tax basis for $790,000 in February. The purchaser paid $83,000 cashRequired ARequired B RequiredC Is this difference favorable or unfavorable? OFavorable OUnfavorable Required C Required ARequired A Required B Required C Compute the difference between TPWs book and tax income resulting from the installme percen
TPW, a calendar year taxpayer, sold land with a $543,000 tax basis for $790,000 in February. The purchaser paid $83,000 cash at closing and gave TPW an interest- bearing note for the $707,000 remaining price. In August, TPW received a $58,550 payment from the purchaser consisting of a $35,350 principal payment and a $23,200 interest payment. Assume that TPW uses the installment sale method of accounting. a. Compute the difference between TPW's book and tax income resulting from the installment sale method. b. Is this difference favorable or unfavorable? c. Using a 21 percent tax rate, compute PTR's deferred tax asset or liability (identify which) resulting from the book/tax difference. Complete this question by entering your answers in the tabs below. Required ARequired B Required C Compute the difference between TPW's book and tax income resulting from the installme percentage to 2 decimal places, and intermediate calculations to the nearest whole dollar Book/tax difference Required B Required A
Required ARequired B RequiredC Is this difference favorable or unfavorable? OFavorable OUnfavorable Required C Required A
Required A Required B Required C Compute the difference between TPW's book and tax income resulting from the installme percentage to 2 decimal places, and intermediate calculations to the nearest whole dollar K Required B Requiredc
Answer #1

Solution:

a)Calculation of the difference between book income & tax income:

Book income = Amount realized on sale of land - adjusted tax basis in hand

Amount realized on sale of land = 83,000 + 707,000 = $790,000

Adjusted tax basis in land = $543,000

Accordingly , Book income = 790,000 - 543,000 = $247,000

Tax income =Cash received on sale of land * Gross profit percent = 118,350 * 31.27% = 37,008

Note :

1)Cash received on sale of land = 83,000 + 35,350 = $118,350

2)Gross profit percent = 247,000 / 790,000 = 31.27%

Difference between book income & tax income = $247,000 - $37,008 = $209,991

b)The difference between book income & tax income is favorable

c)Deferred tax liability = 21% of 209,991 = $44,098